### v3

@@ 2.1.3 Detection Head
The detection head follows the standard YOLO 11 design without modification. In the FPN pathway, fused $P_{5}$ features are upsampled $2\times$ (nearest-neighbor) and concatenated with fused $P_{4}$, then upsampled again and concatenated with RGB-only $P_{3}$; a C3k2 block refines the concatenated features at each level. The PAN pathway reverses this flow via stride-2 convolutions, propagating high-resolution spatial detail from $P_{3}$ back toward $P_{4}$ and $P_{5}$.
Each scale terminates in a decoupled head with separate classification and regression branches. Bounding boxes are predicted via distribution focal loss (DFL), regressing discretized distributions over left-top-right-bottom offsets relative to anchor points. Positive samples are assigned by the Task-Aligned Assigner, which scores candidates jointly by classification confidence and localization quality. The training loss combines binary cross-entropy for classification, CIoU for box regression, and DFL for distribution regression, all with default YOLO 11 weights. The proposed contributions are thus confined entirely to the feature extraction and fusion stages; the detection head and training procedure remain identical to the baseline.
@@ 2.2 Semantic Recomposition Module
SRM is designed to suppress modality-specific noise without eroding the fine-grained details on which small-defect localization depends. In transmission-line imagery, background textures (conductor strand patterns, fitting reflections), structural edges (conductor–sky boundaries), and depth sensor artifacts (holes, edge bleeding, quantization noise) all produce strong feature responses that inflate false-positive rates at the very scales most relevant to small targets. Channel and spatial attention can reweight existing activations but cannot impose structural constraints on what constitutes valid content. SRM takes a different approach: it compresses features into a learned low-dimensional subspace, refines them within that subspace, and reinjects the result through controlled residual mixing.
Given an input feature map $\mathbf{F}\in\mathbb{R}^{B\times C\times H\times W}$, the pipeline proceeds as follows (Fig. 2). The encoding stage projects $\mathbf{F}$ into a $K$-dimensional latent space via a $1\times 1$ pointwise convolution without BN or activation:
@@
The reduction from $C$ to $K$ imposes an information bottleneck that retains only components admitting a stable low-dimensional representation, discarding incidental variation.
@@
No activation function follows this convolution: keeping the refinement linear prevents nonlinearity from coupling noise across latent channels. The $5\times 5$ kernel captures local relationships between adjacent structural elements while maintaining parameter efficiency through the depthwise formulation.
@@
where $\epsilon=10^{-5}$. Unlike batch or layer normalization, which pool statistics over spatial or batch dimensions, position-wise normalization computes statistics at each $(h,w)$ location independently. Transmission-line images exhibit pronounced spatial heterogeneity—sky, metallic towers, and insulator assemblies produce very different activation patterns—and sharing statistics across these regions would suppress informative variation. A consequence of this design is that absolute activation magnitudes are erased at each position; only the relative distribution across latent channels survives. This reduces sensitivity to the systematic scale difference between RGB and depth activations at corresponding locations, but also discards magnitude information that may be relevant to confidence estimation. The residual mixing in the final stage partially compensates by preserving a fraction of the original features.
The decoding stage maps the normalized representation back to $C$ channels via a $1\times 1$ pointwise convolution (without BN or activation):
@@
The output is a convex combination of the refined and original features:
@@
where $\alpha=0.8$ is a fixed scalar. Weighting the refined branch heavily maximizes noise suppression, while the $(1{-}\alpha)$ residual preserves fine-grained spatial detail that the bottleneck cannot fully reconstruct—a balance between purification strength and localization fidelity.
All six instances share the same architecture but serve distinct functional roles via the choice of $K$. At the branch level (K=128K{=}128), the compact bottleneck filters modality-private noise—illumination artifacts in RGB, sensor holes and quantization errors in depth—by compressing features before normalization, producing per-modality representations with narrower inter-modal distribution gaps. At the fusion level (K=256K{=}256), the expanded bottleneck operates on concatenated cross-modal features; position-wise normalization then acts on the joint RGB-depth latent space, attenuating channels whose activations are inconsistent across modalities while reinforcing those that exhibit coherent patterns. In both cases, $K$ is set to roughly half the input channel dimension, balancing compression strength against representational capacity.
@@ 2.3 Contextual Semantic Integration Framework
Because SRM relies on local convolutions ($5{\times}5$ depthwise kernel) and per-position normalization, it has no mechanism for capturing long-range spatial dependencies. At the $P_{5}$ fusion level, however, defect discrimination often hinges on global structural context: a damaged insulator is easier to identify when the network can reference the regular arrangement of intact neighbors, and a bird nest is distinguished from legitimate equipment by its inconsistency with the expected component layout. To provide this capability, CMAFNet inserts CSIF at the $P_{5}$ fusion stage, between C3k2 and the terminal SRM.
CSIF is placed exclusively at $P_{5}$ for two reasons. Global attention is most effective where features encode abstract relational concepts rather than fine-grained spatial patterns, and the quadratic cost of self-attention ($\mathcal{O}(H^{2}W^{2})$) is tractable only at the lowest resolution ($20{\times}20$ at the nano scale); deploying CSIF at $P_{3}$ or $P_{4}$ would multiply FLOPs by $16\times$ or $4\times$ with no commensurate benefit.
@@
As illustrated in Fig. 3, CSIF processes an input $\mathbf{X}\in\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ through a project–refine–project pipeline. A $1{\times}1$ convolutional stem with BN and SiLU expands the channel dimension by a factor of e=2e{=}2, producing a $2C_{\mathrm{in}}$-channel tensor that is then split into two equal halves along the channel axis. One half—the attention portion—passes through a sequence of $n$ Contextual Semantic Interaction Blocks (CSIBs); the other—the bypass portion—is forwarded without modification. After $n$ iterations the two halves are concatenated, and a $1{\times}1$ convolution with BN and SiLU maps the result to the target dimension $C_{\mathrm{out}}$. Restricting attention to half the channels has two benefits. It confines global interaction to a subspace of the representation, mitigating the over-smoothing that full-channel attention would inflict on small-object features, and it halves the computational cost of the attention operation.
@@
Each CSIB (Fig. 4) implements a standard attention–FFN cycle with ASRM replacing conventional layer normalization in a post-normalization arrangement. The first phase computes multi-head self-attention:
@@
with $h=\max(1,\lfloor C/64\rfloor)$ heads and per-head dimension $d_{k}=C/h$. A residual connection yields $\mathbf{U}=\mathbf{H}+\mathrm{Attn}(\mathbf{H})$, followed by ASRM: $\hat{\mathbf{U}}=\mathrm{ASRM}(\mathbf{U})$. The second phase applies a two-layer feed-forward network:
@@
where $\sigma$ is SiLU and the hidden dimension is $2C$. A second residual connection and ASRM produce the block output: $)\mathbf{H}^{(i)}=\mathrm{ASRM}(\hat{\mathbf{U}}+\mathrm{FFN}(\hat{\mathbf{U}}$)). Post-normalization is chosen over pre-normalization because the concatenated RGB-depth features carry persistent distribution heterogeneity; normalizing after the residual addition corrects this shift before the next transformation can amplify it.
ASRM is designed to regulate the distribution heterogeneity of cross-modal fused features more flexibly than standard layer normalization. Given $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, it first computes a gated interpolation between the layer-normalized and original representations:
@@
where $\mathrm{LN}(\cdot)$ is channel-wise layer normalization (statistics computed over $C$ at each spatial position), $\boldsymbol{\gamma},\boldsymbol{\gamma}_{x}\in\mathbb{R}^{C\times 1\times 1}$ are learnable per-channel scales initialized to $\boldsymbol{\gamma}{=}0.01$ and $\boldsymbol{\gamma}_{x}{=}1$, and $\odot$ denotes the Hadamard product. With this initialization ASRM approximates an identity mapping at the outset, avoiding premature interference with the still-unstable attention weights; as training proceeds, the gate gradually shifts toward the normalized branch, effecting a smooth transition from weak to strong regulation.
@@
where $\mathcal{P}_{\mathrm{down}}{:}\;\mathbb{R}^{C}{\to}\mathbb{R}^{r}$ is a $1{\times}1$ convolution reducing channels to r=64r{=}64, $\mathrm{DW}_{3,5,7}$ denotes three parallel depthwise convolutions with kernel sizes $3{\times}3$, $5{\times}5$, and $7{\times}7$ whose outputs are summed, $\sigma$ is GELU, Dropout operates with probability 0.05, and $\mathcal{P}_{\mathrm{up}}{:}\;\mathbb{R}^{r}{\to}\mathbb{R}^{C}$ restores the original channel dimension. The final output is:
@@
Because $r{=}64\ll C$, the residual formulation confines the adaptation to a low-rank subspace, ensuring that $\mathcal{F}$ acts as a correction rather than a full feature rewrite. The multi-scale depthwise convolutions supply spatially adaptive context at three receptive-field sizes without the quadratic cost of a second attention layer. This combination distinguishes ASRM from both standard gated normalization—which lacks a spatial adaptation path—and adapter-style modules—which lack the gated normalization front-end that governs the identity-to-regulation transition.
CSIF uses n=2n{=}2 CSIB blocks at $P_{5}$, balancing modeling capacity against the over-smoothing that additional blocks would introduce at $20{\times}20$ resolution. Through global attention, the module enables direct interaction between RGB-derived and depth-derived features across all spatial positions, letting the network learn cross-modal correspondences beneficial to defect detection. The refined $P_{5}$ features then pass through the terminal SRM for a final round of bottleneck-based distribution normalization before entering the detection head.
@@ 3.1 Datasets
This study employs the TL-RGBD (Transmission Line RGB-D Dataset), a large-scale, high-resolution RGB-D dataset specifically designed for detecting defects in power transmission line components. The dataset comprises 10,000 paired RGB and depth images of insulators, tie wires, and poles, captured under authentic inspection scenarios by China Southern Power Grid using UAVs equipped with synchronized RGB cameras and depth sensors. The depth information, acquired directly through onboard depth cameras during flight missions, provides complementary geometric and spatial information crucial for precise defect localization.
@@
The TL-RGBD dataset exhibits several distinctive characteristics that reflect the intrinsic challenges of UAV-based transmission line inspection. Statistical analysis reveals a highly imbalanced distribution of object sizes, with each annotated object occupying an average area of 422.12 pixels. The dataset composition demonstrates an extreme dominance of small-scale targets: 94.51% are classified as small objects (69,413 instances with area $32^{2}$ pixels), 5.48% as medium-sized objects (4,022 instances with $32^{2}\leq$ area $96^{2}$ pixels), and merely 0.02% as large objects (13 instances with area $\geq 96^{2}$ pixels). This extreme skew toward small-scale targets poses significant challenges for detection algorithms and accurately represents real-world inspection scenarios where defects often manifest as subtle, localized anomalies.
For experimental evaluation, the TL-RGBD dataset was randomly divided into training, validation, and test subsets following an 8:1:1 split ratio. Specifically, 80% of the image pairs were used for model training, 10% for validation and hyperparameter selection, and the remaining 10% were reserved for final performance evaluation. The split was conducted at the image level to ensure that paired RGB and depth images always remain within the same subset. This partitioning strategy ensures sufficient training data while enabling unbiased validation and testing under consistent data distributions.
The depth maps, captured synchronously with RGB images using onboard depth sensors, provide critical geometric cues that complement RGB appearance features. Depth data proves particularly valuable for distinguishing between visually similar defects, identifying component protrusions or deformations, measuring geometric anomalies, and improving localization accuracy under challenging illumination conditions where RGB information alone may be insufficient. This multimodal approach enables more robust defect characterization by leveraging both photometric and geometric properties of transmission line components.
Figure 5: Sample image pairs from TL-RGBD dataset. (a) and (c) are RGB images captured by UAV-mounted cameras showing transmission line components under real inspection conditions. (b) and (d) are the corresponding depth maps acquired synchronously, providing geometric information for defect localization.
@@ 3.2 Model Training and Evaluation Metrics
Given the prevalence of small-scale defects in transmission-line inspection scenarios captured by UAV-mounted RGB-D sensors, we additionally report scale-sensitive metrics defined by the COCO protocol: $\text{AP}_{s}$ for small objects (area $32^{2}$ pixels) and $\text{AP}_{m}$ for medium objects (area between $32^{2}$ and $96^{2}$ pixels). The $\text{AP}_{s}$ metric is particularly informative for assessing the model’s capability to detect fine-grained, low-pixel-count anomalies—a critical requirement for reliable transmission-line monitoring. Computational efficiency and model complexity were quantified using floating-point operations (FLOPs) and the total number of parameters (Params, in millions). Detection performance was evaluated through precision ($P$), recall ($R$), average precision (AP), and mean average precision (mAP), defined as follows:
@@
where $TP$, FPFP, and $FN$ denote true positives, false positives, and false negatives, respectively. For each defect category, average precision is computed as the area under the precision-recall curve:

### v4

@@ 2.1.3 Detection Head
The detection head follows the standard YOLO 11 design without modification. In the FPN pathway, fused $P_{5}$ features are upsampled by a factor of $2\times$ (nearest-neighbor) and concatenated with fused $P_{4}$, then upsampled again and concatenated with RGB-only $P_{3}$; a C3k2 block refines the concatenated features at each level. The PAN pathway reverses this flow through stride-2 convolutions, propagating high-resolution spatial detail from $P_{3}$ back toward $P_{4}$ and $P_{5}$.
Each scale terminates in a decoupled head with separate classification and regression branches. Bounding boxes are predicted through distribution focal loss (DFL), which regresses discretized distributions over left-top-right-bottom offsets relative to anchor points. Positive samples are assigned by the Task-Aligned Assigner, which scores candidates jointly on the basis of classification confidence and localization quality. The training loss combines binary cross-entropy for classification, CIoU for box regression, and DFL for distribution regression, all with default YOLO 11 weights. The proposed contributions are thus confined entirely to the feature extraction and fusion stages; the detection head and training procedure remain identical to those of the baseline.
@@ 2.2 Semantic Recomposition Module
SRM is designed to suppress modality-specific noise without eroding the fine-grained details on which small-defect localization depends. In transmission-line imagery, background textures such as conductor strand patterns and fitting reflections, structural edges at conductor–sky boundaries, and depth sensor artifacts including holes, edge bleeding, and quantization noise all produce strong feature responses that inflate false-positive rates at the scales most relevant to small targets. Channel and spatial attention mechanisms can reweight existing activations but cannot impose structural constraints on what constitutes valid content. SRM adopts a fundamentally different strategy: it compresses features into a learned low-dimensional subspace, refines them within that subspace, and reinjects the result through controlled residual mixing.
Given an input feature map $\mathbf{F}\in\mathbb{R}^{B\times C\times H\times W}$, the pipeline proceeds as follows (Fig. 2). The encoding stage projects $\mathbf{F}$ into a $K$-dimensional latent space via a $1\times 1$ pointwise convolution without batch normalization or activation:
@@
The reduction from $C$ to $K$ introduces an information bottleneck that retains only components admitting a stable low-dimensional representation while discarding incidental variation.
@@
No activation function follows this convolution; maintaining linearity in the refinement stage prevents nonlinear coupling of noise across latent channels. The $5\times 5$ kernel is sufficiently large to capture local relationships among adjacent structural elements while preserving parameter efficiency through the depthwise formulation.
@@
where $\epsilon=10^{-5}$. Unlike batch or layer normalization, which pool statistics over spatial or batch dimensions, position-wise normalization computes statistics at each $(h,w)$ location independently. Transmission-line images exhibit pronounced spatial heterogeneity—sky regions, metallic towers, and insulator assemblies produce markedly different activation patterns—and sharing statistics across these regions would suppress informative variation. A consequence of this design is that absolute activation magnitudes are erased at each position, leaving only the relative distribution across latent channels. This property reduces sensitivity to the systematic scale discrepancy between RGB and depth activations at corresponding locations, yet it also discards magnitude information that may be relevant to confidence estimation. The residual mixing in the final stage partially compensates for this loss by preserving a fraction of the original features.
The decoding stage maps the normalized representation back to $C$ channels via a $1\times 1$ pointwise convolution without batch normalization or activation:
@@
The output is formed as a convex combination of the refined and original features:
@@
where $\alpha=0.8$ is a fixed scalar. Assigning a large weight to the refined branch maximizes noise suppression, while the $(1{-}\alpha)$ residual preserves fine-grained spatial detail that the bottleneck cannot fully reconstruct, thereby striking a balance between purification strength and localization fidelity.
All six instances share the same architecture but serve distinct functional roles determined by the choice of $K$. At the branch level (K=128K{=}128), the compact bottleneck filters modality-private noise—illumination artifacts in RGB and sensor holes together with quantization errors in depth—by compressing features before normalization, yielding per-modality representations with narrower inter-modal distribution gaps. At the fusion level (K=256K{=}256), the expanded bottleneck operates on concatenated cross-modal features; position-wise normalization then acts on the joint RGB-depth latent space, attenuating channels whose activations are inconsistent across modalities while reinforcing those that exhibit coherent patterns. In both cases, $K$ is set to approximately half the input channel dimension, balancing compression strength against representational capacity.
@@ 2.3 Contextual Semantic Integration Framework
Because SRM relies on local convolutions ($5{\times}5$ depthwise kernel) and per-position normalization, it lacks the capacity to capture long-range spatial dependencies. At the $P_{5}$ fusion level, however, defect discrimination often depends on global structural context: a damaged insulator is easier to identify when the network can reference the regular arrangement of intact neighbors, and a bird nest is distinguished from legitimate equipment by its inconsistency with the expected component layout. To provide this capability, CMAFNet inserts CSIF at the $P_{5}$ fusion stage, between C3k2 and the terminal SRM.
CSIF is placed exclusively at $P_{5}$ for two reasons. First, global attention is most effective when features encode abstract relational semantics rather than fine-grained spatial patterns. Second, the quadratic cost of self-attention ($\mathcal{O}(H^{2}W^{2})$) is tractable only at the lowest resolution ($20{\times}20$ at the nano scale); deploying CSIF at $P_{3}$ or $P_{4}$ would multiply FLOPs by $16\times$ or $4\times$, respectively, with no commensurate gain.
@@
As illustrated in Fig. 3, CSIF processes an input $\mathbf{X}\in\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ through a project–refine–project pipeline. A $1{\times}1$ convolutional stem with BN and SiLU expands the channel dimension by a factor of e=2e{=}2, producing a $2C_{\mathrm{in}}$-channel tensor that is then split into two equal halves along the channel axis. One half, referred to as the attention portion, passes through a sequence of $n$ Contextual Semantic Interaction Blocks (CSIBs), while the other, referred to as the bypass portion, is forwarded without modification. After $n$ iterations, the two halves are concatenated, and a $1{\times}1$ convolution with BN and SiLU maps the result to the target dimension $C_{\mathrm{out}}$. Restricting attention to half the channels offers two advantages: it confines global interaction to a subspace of the representation, thereby mitigating the over-smoothing that full-channel attention would impose on small-object features, and it halves the computational cost of the attention operation.
@@
Each CSIB (Fig. 4) implements a standard attention–FFN cycle, with ASRM replacing conventional layer normalization in a post-normalization arrangement. The first phase computes multi-head self-attention:
@@
with $h=\max(1,\lfloor C/64\rfloor)$ heads and a per-head dimension of $d_{k}=C/h$. A residual connection yields $\mathbf{U}=\mathbf{H}+\mathrm{Attn}(\mathbf{H})$, followed by ASRM: $\hat{\mathbf{U}}=\mathrm{ASRM}(\mathbf{U})$. The second phase applies a two-layer feed-forward network:
@@
where $\sigma$ denotes SiLU and the hidden dimension is $2C$. A second residual connection and ASRM produce the block output: $)\mathbf{H}^{(i)}=\mathrm{ASRM}(\hat{\mathbf{U}}+\mathrm{FFN}(\hat{\mathbf{U}}$)). Post-normalization is preferred over pre-normalization because the concatenated RGB-depth features exhibit persistent distribution heterogeneity; normalizing after the residual addition corrects this distributional shift before the subsequent transformation can amplify it.
ASRM is designed to regulate the distribution heterogeneity of cross-modal fused features more flexibly than standard layer normalization. Given $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, it first computes a gated interpolation between the layer-normalized representation and the original input:
@@
where $\mathrm{LN}(\cdot)$ denotes channel-wise layer normalization with statistics computed over $C$ at each spatial position, $\boldsymbol{\gamma},\boldsymbol{\gamma}_{x}\in\mathbb{R}^{C\times 1\times 1}$ are learnable per-channel scales initialized to $\boldsymbol{\gamma}{=}0.01$ and $\boldsymbol{\gamma}_{x}{=}1$, and $\odot$ denotes the Hadamard product. Under this initialization, ASRM approximates an identity mapping at the outset, thereby avoiding premature interference with the still-unstable attention weights. As training proceeds, the gate gradually shifts toward the normalized branch, effecting a smooth transition from weak to strong regulation.
@@
where $\mathcal{P}_{\mathrm{down}}{:}\;\mathbb{R}^{C}{\to}\mathbb{R}^{r}$ is a $1{\times}1$ convolution that reduces channels to r=64r{=}64, $\mathrm{DW}_{3,5,7}$ denotes three parallel depthwise convolutions with kernel sizes of $3{\times}3$, $5{\times}5$, and $7{\times}7$ whose outputs are summed, $\sigma$ is GELU, Dropout operates with a probability of 0.05, and $\mathcal{P}_{\mathrm{up}}{:}\;\mathbb{R}^{r}{\to}\mathbb{R}^{C}$ restores the original channel dimension. The final output is:
@@
Because $r{=}64\ll C$, the residual formulation confines the adaptation to a low-rank subspace, ensuring that $\mathcal{F}$ acts as a correction rather than a complete feature rewrite. The multi-scale depthwise convolutions provide spatially adaptive context at three receptive-field sizes without incurring the quadratic cost of a second attention layer. This combination distinguishes ASRM from both standard gated normalization, which lacks a spatial adaptation path, and adapter-style modules, which lack the gated normalization front-end that governs the identity-to-regulation transition.
CSIF uses n=2n{=}2 CSIB blocks at $P_{5}$, striking a balance between modeling capacity and the over-smoothing that additional blocks would introduce at $20{\times}20$ resolution. Through global attention, the module enables direct interaction between RGB-derived and depth-derived features across all spatial positions, allowing the network to learn cross-modal correspondences that are beneficial for defect detection. The refined $P_{5}$ features then pass through the terminal SRM for a final round of bottleneck-based distribution normalization before entering the detection head.
@@ 3.1 Datasets
This study employs the TL-RGBD (Transmission Line RGB-D Dataset), a large-scale, high-resolution RGB-D dataset specifically designed for detecting defects in power transmission line components. The dataset comprises 10,000 paired RGB and depth images of insulators, tie wires, and poles, captured under authentic inspection scenarios by China Southern Power Grid using UAVs equipped with synchronized RGB cameras and depth sensors. The depth information, acquired directly through onboard depth cameras during flight missions, provides complementary geometric and spatial cues essential for precise defect localization.
@@
The TL-RGBD dataset exhibits several distinctive characteristics that reflect the intrinsic challenges of UAV-based transmission line inspection. Statistical analysis reveals a highly imbalanced distribution of object sizes, with each annotated instance occupying an average area of 422.12 pixels. The dataset composition is dominated by small-scale targets: 94.51% of instances are classified as small objects (69,413 instances with area $32^{2}$ pixels), 5.48% as medium-sized objects (4,022 instances with $32^{2}\leq$ area $96^{2}$ pixels), and merely 0.02% as large objects (13 instances with area $\geq 96^{2}$ pixels). This extreme skew toward small-scale targets poses significant challenges for detection algorithms and faithfully represents real-world inspection scenarios where defects often manifest as subtle, localized anomalies.
For experimental evaluation, the TL-RGBD dataset was randomly divided into training, validation, and test subsets following an 8:1:1 split ratio. 80% of the image pairs were used for model training, 10% for validation and hyperparameter selection, and the remaining 10% were reserved for final performance evaluation. The split was conducted at the image level to ensure that paired RGB and depth images always reside within the same subset. This partitioning strategy provides sufficient training data while enabling unbiased validation and testing under consistent data distributions.
The depth maps, captured synchronously with RGB images using onboard depth sensors, encode critical geometric cues that complement RGB appearance features. Depth data is particularly valuable for distinguishing between visually similar defects, identifying component protrusions or deformations, quantifying geometric anomalies, and improving localization accuracy under challenging illumination conditions where RGB information alone may be insufficient. This multimodal approach enables more robust defect characterization by leveraging both photometric and geometric properties of transmission line components.
Figure 5: Sample image pairs from the TL-RGBD dataset. (a) and (c) are RGB images captured by UAV-mounted cameras showing transmission line components under real inspection conditions. (b) and (d) are the corresponding depth maps acquired synchronously, providing geometric information for defect localization.
@@ 3.2 Model Training and Evaluation Metrics
Given the prevalence of small-scale defects in transmission-line inspection scenarios captured by UAV-mounted RGB-D sensors, we additionally report scale-sensitive metrics defined by the COCO protocol: $\text{AP}_{s}$ for small objects (area $32^{2}$ pixels) and $\text{AP}_{m}$ for medium objects (area between $32^{2}$ and $96^{2}$ pixels). The $\text{AP}_{s}$ metric is particularly informative for assessing the capability of a model to detect fine-grained, low-pixel-count anomalies—a critical requirement for reliable transmission-line monitoring. Computational efficiency and model complexity were quantified using floating-point operations (FLOPs) and the total number of parameters (Params, in millions). Detection performance was evaluated through precision ($P$), recall ($R$), average precision (AP), and mean average precision (mAP), defined as follows:
@@
where $TP$, FPFP, and $FN$ denote true positives, false positives, and false negatives, respectively. For each defect category, the average precision is computed as the area under the precision-recall curve: